OpenStudy (anonymous):

solve the radical equation √18x-9=x+4

OpenStudy (anonymous):

there are a few possible first steps.. what do you think?

OpenStudy (anonymous):

square it.

OpenStudy (anonymous):

yea, then it's: 18x-9 = x^2 +8x +16 now what?

OpenStudy (anonymous):

add 9

OpenStudy (anonymous):

ok, then it's: 18x = x^2 +8x +25 now what?

OpenStudy (anonymous):

subtract 8x?

OpenStudy (anonymous):

how is it 0?

OpenStudy (anonymous):

opps. lol

OpenStudy (anonymous):

lets subtrack 18x... then it's x^2 -10x + 25

OpenStudy (anonymous):

=0

OpenStudy (anonymous):

Okay, now factor it ?

OpenStudy (anonymous):

totally

OpenStudy (anonymous):

(x-5)(x-5)

OpenStudy (anonymous):

so my answers 5?

OpenStudy (anonymous):

so x = 5

OpenStudy (anonymous):

idneeeeed.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

sure!